# Spectral and Temporal Denoising for Differentially Private Optimization

### v1

@@ Abstract
This paper introduces the FFT-Enhanced Kalman Filter (FFTKF), a differentially private optimization method that addresses the challenge of preserving performance in DP-SGD, where added noise typically degrades model utility. FFTKF integrates frequency-domain noise shaping with Kalman filtering to enhance gradient quality while preserving $(\varepsilon,\delta)$-DP guarantees. It employs a high-frequency shaping mask in the Fourier domain to concentrate differential privacy noise in less informative spectral components, preserving low-frequency gradient signals. A scalar-gain Kalman filter with finite-difference Hessian approximation further refines the denoised gradients. With a per-iteration complexity of $\mathcal{O}(d\log d)$, FFTKF demonstrates improved test accuracy over DP-SGD and DiSK across MNIST, CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets using CNNs, Wide ResNets, and Vision Transformers. Theoretical analysis confirms that FFTKF maintains equivalent privacy guarantees while achieving a tighter privacy-utility trade-off through reduced noise and controlled bias.
@@ 1 Introduction
Differential Privacy (DP) has become a foundational concept for safeguarding individual-level information in machine learning and data analysis, providing rigorous guarantees against information leakage from model outputs [1, 30, 25, 18]. Standard DP mechanisms, such as the Laplace and Gaussian mechanisms, achieve privacy by injecting calibrated noise into data or gradients. However, this noise injection often results in substantial degradation in model utility, particularly in high-dimensional or deep learning scenarios.
One of the central challenges in differentially private learning is improving the performance of DP-SGD, a widely used algorithm that combines stochastic gradient descent with Gaussian noise perturbation [1]. Due to the stochastic and high-magnitude nature of DP noise, especially under tight privacy budgets, the privatized gradients tend to exhibit poor signal-to-noise ratios. Consequently, effectively denoising DP-SGD updates to enhance performance while strictly preserving DP guarantees remains a formidable and open problem. Recent studies highlight that the noise introduced in DP-SGD can significantly impair convergence and model accuracy, particularly in complex tasks, making denoising a critical yet challenging objective [23, 26]. These works underscore the difficulty of balancing noise reduction with the preservation of privacy constraints, as excessive denoising risks compromising the $(\epsilon,\delta)$-DP guarantees.
To address this, recent methods have introduced signal processing and state estimation tools to enhance gradient quality under DP. For example, the DiSK framework [32] incorporates Kalman filtering into DP optimization to recursively estimate cleaner gradients, resulting in improved convergence and test accuracy. Kalman filters are well-established tools for estimating latent states in dynamic systems based on noisy observations [19, 20]. Their application in DP optimization allows for smoother updates by exploiting temporal correlations across gradient steps, which is particularly effective in large-scale and deep models.
In parallel, frequency-domain techniques such as low-pass filtering have shown promise in separating useful signal components from high-frequency noise [9, 17, 22]. Leveraging the Fast Fourier Transform (FFT), recent signal processing approaches selectively attenuate noise by filtering out high-frequency components in a computationally efficient manner [6]. When adapted to machine learning, these spectral methods provide a structured way to shape and redistribute noise, preserving signal fidelity in lower-frequency domains [2, 26].
Building on these developments, we propose the FFT-Enhanced Kalman Filter (FFTKF), a novel optimization framework that combines spectral noise shaping via FFT with Kalman filtering for improved gradient estimation under differential privacy. Our method reshapes the injected DP noise to concentrate its energy in high-frequency spectral components, thereby preserving low-frequency gradient information that is critical for effective model training. The resulting privatized gradients are then denoised using a scalar-gain Kalman filter that leverages temporal structure to produce stable update directions.
Our contributions are as follows. First, we introduce a frequency-domain noise shaping strategy that integrates into the DP optimization pipeline while preserving the required $(\epsilon,\delta)$-DP guarantees. Second, we combine this with a simplified Kalman filter that enables efficient gradient tracking with a per-step complexity of $O(d\log d)$. Third, we empirically validate FFTKF on standard benchmarks—MNIST, CIFAR-10, CIFAR-100, and Tiny-ImageNet—across multiple model architectures including CNNs, Wide ResNets, and Vision Transformers. Our results show that FFTKF consistently improves test accuracy over DP-SGD and DiSK, particularly under tight privacy budgets. These findings demonstrate the effectiveness of combining spectral filtering and state estimation for denoising gradients in private optimization.
@@ 2 Related Works
Gradient-Based Optimization. Stochastic Gradient Descent (SGD) and its variants, such as Adam, form the backbone of optimization in machine learning by iteratively updating model parameters using gradient information [24, 12]. SGD relies on noisy gradient estimates computed from random mini-batches, offering computational efficiency but suffering from variance in updates. Adam enhances SGD by incorporating adaptive momentum and second-moment estimates, which accelerate convergence, particularly in non-convex settings. However, these methods lack inherent privacy protections, as gradients can leak sensitive information about training data. This limitation has spurred the development of privacy-preserving optimizers like DP-SGD, which integrate noise to safeguard data privacy [1].
Differential Privacy Optimization. Differential Privacy (DP) is a rigorous framework for protecting individual data privacy in machine learning by injecting calibrated noise into data or gradients, often at the cost of reduced model utility [1, 30, 25, 18]. DP-SGD, a widely adopted algorithm, achieves privacy by adding Gaussian noise to gradients during stochastic gradient descent, but the high-magnitude noise often degrades signal-to-noise ratios, impairing convergence [1]. Recent efforts address this trade-off through innovative techniques. For instance, Fan and Xiong proposed an adaptive noise adjustment method for real-time aggregate monitoring, dynamically balancing privacy and utility based on data characteristics [10]. Similarly, Zhang et al.’s DiSK framework integrates Kalman filtering to denoise privatized gradients, enhancing model performance while adhering to privacy constraints [32]. Adaptive clipping techniques further improve DP-SGD by dynamically tuning noise levels to preserve gradient information [28].
Kalman Filter Optimization. Kalman filters are powerful tools for state estimation in noisy dynamic systems, leveraging temporal correlations to refine estimates from noisy observations [19, 20]. In the context of differential privacy, the DiSK framework employs a simplified Kalman filter to smooth noisy gradient estimates, improving convergence rates and reducing computational overhead in DP optimization [32]. This approach is particularly effective for large-scale models, where noise can significantly disrupt training. Recent research has also explored Kalman filters in federated learning, where they help balance privacy and accuracy in distributed environments by refining gradient updates across clients. These advancements highlight the versatility of Kalman-based methods in privacy-preserving machine learning.
Low-Pass Filtering Techniques. Low-pass filters are essential in signal processing for attenuating high-frequency noise while preserving critical signal components, making them valuable for data analysis [6]. Fourier transform-based low-pass filters, known for their computational efficiency, are widely used to separate noise from meaningful signals [3]. In differential privacy, adaptive low-pass filtering techniques optimize cutoff frequencies to maximize utility under strict privacy budgets, effectively denoising gradients in high-dimensional tasks [2, 26, 5]. These methods are particularly relevant for deep learning, where preserving low-frequency gradient information is crucial for maintaining model performance while ensuring robust privacy guarantees.
@@ 3 Methodology
Figure 1: Visualization of the proposed frequency-domain gradient denoising process. (A) The original flattened gradient before privatization. (B) The FFT transforms the gradient into the frequency domain, where shaped Gaussian noise is added and attenuated by a high-frequency mask. Kalman filtering is then applied in the time domain to exploit temporal correlations. (C) The final denoised gradient (dashed line) shows reduced high-frequency perturbations while preserving the underlying signal structure.
Preliminaries. Let $\mathbb{R}^{d}$ denote the $d$-dimensional Euclidean space. A vector $z\in\mathbb{R}^{d}$ is written as $z=(z_{0},\dots,z_{d-1})^{\top}$, with $\ell_{2}$-norm $\|z\|_{2}=\sqrt{\sum_{i=0}^{d-1}z_{i}^{2}}$. The identity matrix in $\mathbb{R}^{d\times d}$ is $I_{d}$, and a diagonal matrix with entries $\varphi_{0},\dots,\varphi_{d-1}$ is $\mathrm{diag}(\varphi_{0},\dots,\varphi_{d-1})$. For a function $f:\mathbb{R}^{d}\to\mathbb{R}$, its gradient is $\nabla f$, and Hessian is $\nabla^{2}f$. The Hadamard product is $\odot$, and $\lfloor x\rfloor$ denotes the floor function. The Gaussian distribution with mean zero and covariance $\sigma^{2}I_{d}$ is $\mathcal{N}(0,\sigma^{2}I_{d})$. We aim to minimize $F(x)=\mathbb{E}_{\xi\sim\mathcal{D}}[f(x;\xi)]$, the expected loss over a data distribution $\mathcal{D}$. Parameters at iteration $t$ are $x_{t}\in\mathbb{R}^{d}$, updated as $x_{t+1}=x_{t}-\eta\tilde{g}_{t}$, where $\eta>0$ is the learning rate and $\tilde{g}_{t}$ is the estimated gradient. The true gradient $\nabla F(x_{t})$ is approximated via a mini-batch $\mathcal{B}_{t}\subseteq\mathcal{D}$ of size $B$, giving $g_{t}=\frac{1}{B}\sum_{\xi\in\mathcal{B}_{t}}\nabla f(x_{t};\xi)$. The parameter difference is $d_{t}=x_{t+1}-x_{t}$. In differential privacy, a mechanism satisfies $(\varepsilon,\delta)$-DP if, for neighboring datasets $D,D^{\prime}$ differing in one entry, and any output set $S$, $\Pr[\mathcal{M}(D)\in S]\leq e^{\varepsilon}\Pr[\mathcal{M}(D^{\prime})\in S]+\delta$. Gradients are clipped as $\text{clip}(v,C)=v\cdot\min(1,C/\|v\|_{2})$ to bound their $\ell_{2}$-norm by $C$, and noise $w_{t}\sim\mathcal{N}(0,\sigma_{w}^{2}I_{d})$ is added.
@@ 3.1 Fast Fourier Transform
This section briefly reviews the discrete Fast Fourier transform (FFT) and the algorithmic considerations that motivate its use for gradient denoising.
Recall that the discrete Fourier transform (DFT) of a real-valued vector $z=(z_{0},\dots,z_{d-1})^{\top}\in\mathbb{R}^{d}$ is the complex vector $\hat{z}=\mathcal{F}(z)\in\mathbb{C}^{d}$, with components
$\hat{z}_{k}=\sum_{n=0}^{d-1}z_{n}\,e^{-2\pi ikn/d},\quad\text{for }k=0,\dots,d-1,$
and its inverse is given by
$z_{n}=\frac{1}{d}\sum_{k=0}^{d-1}\hat{z}_{k}\,e^{2\pi ikn/d},\quad\text{for }n=0,\dots,d-1.$
With this normalization, the Fourier transform $\mathcal{F}$ is unitary, i.e.,
$\mathcal{F}^{-1}(\mathcal{F}(z))=z,$
and Parseval’s identity holds:
$\|z\|_{2}^{2}=\frac{1}{d}\|\hat{z}\|_{2}^{2},$
@@
Consequently, adding Gaussian noise in the Fourier domain preserves the $\ell_{2}$-sensitivity required for $(\varepsilon,\delta)$-DP due to the invariance of the $\ell_{2}$-norm under unitary transformations.
@@
Spectral filtering. A diagonal mask $\Phi=\mathrm{diag}(\varphi_{0},\dots,\varphi_{d-1})$ defines a linear filter
$\mathcal{G}_{\Phi}(z)=\mathcal{F}^{-1}\left(\Phi\,\hat{z}\right)=\frac{1}{d}\sum_{k=0}^{d-1}\varphi_{k}\hat{z}_{k}\,e^{2\pi ikn/d},$
where $\hat{z}=\mathcal{F}(z)$.
@@
$\varphi_{k}=\begin{cases}1,&k<k_{0},\\[3.0pt] 1-\rho\,&k\geq k_{0},\end{cases}$
@@
Remarks. Let $\Phi_{\rho}=\mathrm{diag}(\varphi_{0},\dots,\varphi_{d-1})$, then the filtered version of a privatized gradient $g=\nabla f+w$ is
[…]
When $w\sim\mathcal{N}(0,\sigma^{2}I)$, the transformed noise $\hat{w}:=\mathcal{F}^{-1}(\Phi_{\rho}\cal{F}(w))$ is still zero-mean but now has reduced energy in the low-frequency components:
$\mathbb{E}[\|\hat{w}_{<k_{0}}\|_{2}^{2}]\ll\mathbb{E}[\|\hat{w}\|_{2}^{2}],$
@@ 3.2 Gradient Dynamics with High-Frequency Differential Privacy
$\nabla F(x_{t})=\nabla F(x_{t-1})+\boldsymbol{H}_{t}\!\cdot\!(x_{t}-x_{t-1})+v_{t},$ (1)
where $\boldsymbol{H}_{t}:=\nabla^{2}F(x_{t-1})\in\mathbb{R}^{d\times d}$ is the local Hessian, and $v_{t}$ represents the Taylor expansion remainder:
[…]
Here, $(\cdot)^{\otimes 2}$ denotes the second-order tensor product, capturing the nonlinearity of the third derivative in the remainder [32, 21].
The observed gradient $g_{t}$ is a noisy, privatized estimate of the true gradient:
$g_{t}=\frac{1}{B}\sum_{\xi\in\mathcal{B}_{t}}\text{clip}(\nabla f(x_{t},\xi),C)+w_{t}=C_{t}\nabla F(x_{t})+w^{\prime}_{t},$ (2)
where $w^{\prime}_{t}$ contains both DP noise and subsampling noise. The matrix $C_{t}$ is the effective observation operator determined by the clipping factor and the sampling mechanism. In the ideal full-batch, unclipped setting, $C_{t}=I_{d}$, but more generally it is a contraction mapping satisfying $\|C_{t}\|_{2}\leq 1$.
Combining the update and observation equations, we obtain the system:
@@
To enforce differential privacy while retaining useful structure, we inject noise in the frequency domain:
$g_{t}\;=\;C_{t}\nabla F(x_{t})\;+\;\tilde{w}_{t},\qquad\tilde{w}_{t}:=\mathcal{F}^{-1}\!\bigl{(}\Phi_{\rho}\odot\mathcal{F}(w_{t})\bigr{)},$ (3)
where $w_{t}\sim\mathcal{N}(0,\sigma_{w}^{2}I_{d})$ is isotropic Gaussian noise. The mask $\Phi_{\rho}\in\mathbb{R}^{d}$ is diagonal and satisfies
$(\Phi_{\rho})_{k}=\begin{cases}1,&0\leq k<k_{0},\\[3.0pt] 1-\rho e^{-\alpha(k-k_{0})},&k_{0}\leq k<d,\end{cases}$
for some pivot frequency $k_{0}=\lfloor\lambda d\rfloor$, with shaping strength $\rho\in(0,1)$ and decay rate $\alpha>0$. This ensures that the privacy-preserving noise $\tilde{w}_{t}$ is spectrally shaped to occupy primarily high-frequency components, which contribute less to gradient descent. This leads to improved recoverability of the informative low-frequency gradient content.

### v2

# Fast Fourier Transform-Based Spectral and Temporal Gradient Filtering for Differential Privacy

@@ Abstract
Differential Privacy (DP) has emerged as a key framework for protecting sensitive data in machine learning, but standard DP-SGD often suffers from significant accuracy loss due to injected noise. To address this limitation, we introduce the FFT-Enhanced Kalman Filter (FFTKF), a differentially private optimization method that improves gradient quality while preserving $(\varepsilon,\delta)$-DP guarantees. FFTKF applies frequency-domain filtering to shift privacy noise into less informative high-frequency components, preserving the low-frequency gradient signals that carry most learning information. A scalar-gain Kalman filter with a finite-difference Hessian approximation further refines the denoised gradients. The method has per-iteration complexity $\mathcal{O}(d\log d)$ and achieves higher test accuracy than DP-SGD and DiSK on MNIST, CIFAR-10, CIFAR-100, and Tiny-ImageNet with CNNs, Wide ResNets, and Vision Transformers. Theoretical analysis shows that FFTKF ensures equivalent privacy while delivering a stronger privacy–utility trade-off through reduced variance and controlled bias.
@@ 1 Introduction
Differential Privacy (DP) has become a foundational framework for safeguarding individual-level information in machine learning.
Figure 1: Illustration of the Differentially Private Stochastic Gradient Descent (DP-SGD) process. (A) Original gradients (B) DP-Gradients
This provides rigorous guarantees against information leakage from model outputs [1, 30, 25, 19].
@@
Standard DP mechanisms, such as the Laplace and Gaussian mechanisms, achieve privacy by injecting calibrated noise into data or gradients, as shown in Figure 1. However, this noise often causes significant degradation in model utility, especially in high-dimensional or deep models.
A central challenge in DP learning is improving the performance of DP-SGD [1]. The high variance of DP noise under tight privacy budgets leads to poor signal-to-noise ratios, slowing convergence and reducing accuracy [23, 26]. Thus, denoising while preserving $(\epsilon,\delta)$-DP remains an open problem.
To address this, recent works integrate signal processing and state estimation into DP optimization. The DiSK framework [32] applies Kalman filtering to iteratively estimate cleaner gradients, leveraging temporal correlations [20, 21]. In parallel, frequency-domain methods use low-pass filtering to separate useful gradient signals from high-frequency noise [9, 18, 22, 6, 2, 26]. These advances suggest that combining temporal and spectral denoising can substantially improve the utility of privatized gradients.
Building on this, we propose the FFT-Enhanced Kalman Filter (FFTKF), which reshapes DP noise into high-frequency components via Fast Fourier Transform (FFT) and then applies a scalar-gain Kalman filter to recover stable low-frequency gradients. This approach preserves $(\epsilon,\delta)$-DP while improving convergence and test accuracy.
Our contributions are summarized as follows:
  • A frequency-domain noise shaping strategy that retains DP guarantees.
  • A lightweight Kalman filter update with per-step complexity $O(d\log d)$.
  • Empirical validation on MNIST, CIFAR-10, CIFAR-100, and Tiny-ImageNet across CNNs, Wide ResNets, and Vision Transformers, showing consistent gains over DP-SGD and DiSK.
@@ 2 Related Works
Stochastic Gradient Descent (SGD) and its variants such as Adam are the backbone of modern optimization [24, 13]. SGD provides efficiency by using mini-batch gradients but suffers from high variance, while Adam improves stability through momentum and adaptive scaling. Despite their success, these methods offer no inherent privacy, as gradients may expose sensitive data. This motivated the development of privacy-preserving optimizers such as DP-SGD [1].
Differential Privacy (DP) ensures rigorous protection by injecting calibrated noise into data or gradients [1, 30, 25, 19]. DP-SGD achieves $(\epsilon,\delta)$-DP through Gaussian perturbation but often reduces utility under tight budgets [1]. To mitigate this, adaptive noise adjustment has been explored [10], and the DiSK framework introduced Kalman filtering for denoising [32]. Adaptive clipping further improves learning by tuning gradient norms dynamically [28]. Together, these works highlight the challenge of balancing privacy and accuracy.
Kalman filters estimate hidden states in noisy systems by leveraging temporal correlations [20, 21]. In DP optimization, DiSK applies a simplified Kalman filter to stabilize noisy gradient updates, improving convergence with low computational cost [32]. This temporal smoothing has proven useful in large-scale models and has also been extended to federated learning, where client updates require both privacy and accuracy. These studies show that Kalman-based methods are flexible tools for gradient denoising under DP constraints. Low-pass filters suppress high-frequency noise while preserving dominant low-frequency signals [6, 3]. Fourier-based approaches are especially attractive due to their computational efficiency. In DP, adaptive low-pass filtering has been proposed to maximize utility while meeting privacy budgets, effectively recovering gradient information [2, 26, 5]. These techniques are particularly relevant in deep learning, where most useful gradient information lies in low-frequency components.
@@ 3 Methodology
Preliminaries. We work in $\mathbb{R}^{d}$ with $\ell_{2}$–norm $\|\cdot\|_{2}$. Let $I_{d}$ be the identity and $\mathrm{diag}(\varphi_{0},\ldots,\varphi_{d-1})$ a diagonal matrix. For $f:\mathbb{R}^{d}\!\to\!\mathbb{R}$, denote gradient and Hessian by $\nabla f$ and $\nabla^{2}f$. We use the Hadamard product $\odot$, the floor $\lfloor\cdot\rfloor$, and the Gaussian law $\mathcal{N}(0,\sigma^{2}I_{d})$. We minimize the population loss $F(x)=\mathbb{E}_{\xi\sim\mathcal{D}}[f(x;\xi)]$ via iterates $x_{t+1}=x_{t}-\eta\,\tilde{g}_{t}$ with step size $\eta>0$. Given a mini-batch $\mathcal{B}_{t}$ of size $B$, the stochastic gradient is $g_{t}=\tfrac{1}{B}\sum_{\xi\in\mathcal{B}_{t}}\nabla f(x_{t};\xi)$, and we write the parameter difference as $d_{t}=x_{t+1}-x_{t}$. A mechanism $\mathcal{M}$ is $(\varepsilon,\delta)$-DP if for any neighboring datasets $D,D^{\prime}$ and event $S$, $\Pr[\mathcal{M}(D)\in S]\leq e^{\varepsilon}\Pr[\mathcal{M}(D^{\prime})\in S]+\delta$. In DP-SGD, per-sample gradients are clipped $\mathrm{clip}(v,C)=v\cdot\min\!\bigl{(}1,\,C/\|v\|_{2}\bigr{)}$ and Gaussian noise $w_{t}\!\sim\!\mathcal{N}(0,\sigma_{w}^{2}I_{d})$ is added to enforce privacy.
Figure 2: Visualization of the proposed frequency-domain gradient denoising process. (A) Original gradients before privatization. (B) Differentially private gradients obtained by clipping and adding Gaussian noise. (C) Our FFT+Kalman filtering method denoises the privatized gradients in the frequency domain, reducing high-frequency perturbations while preserving the underlying signal structure.
@@ 3.1 Fast Fourier Transform
This section briefly reviews the discrete Fast Fourier transform (FFT) and the algorithmic considerations that motivate its use for gradient denoising. Recall that the discrete Fourier transform (DFT) of a real-valued vector $z=(z_{0},\dots,z_{d-1})^{\top}\in\mathbb{R}^{d}$ is the complex vector $\hat{z}=\mathcal{F}(z)\in\mathbb{C}^{d}$, with components $\hat{z}_{k}=\sum_{n=0}^{d-1}z_{n}\,e^{-2\pi ikn/d}$ when $k=0,\dots,d-1$ and its inverse is:
$z_{n}$ $=\frac{1}{d}\sum_{k=0}^{d-1}\hat{z}_{k}\,e^{2\pi ikn/d},\quad n=0,\dots,d-1.$ (1)
With this normalization, the Fourier transform $\mathcal{F}$ is unitary is $\mathcal{F}^{-1}(\mathcal{F}(z))=z$ and Parseval’s identity holds:
$\|z\|_{2}^{2}$ $=z^{\ast}z$ (2)
$=z^{\ast}\,\mathcal{F}^{-1}\mathcal{F}\,z$ (3)
$=(\mathcal{F}z)^{\ast}\,(\mathcal{F}z)\cdot\frac{1}{d}$ (4)
$=\frac{1}{d}\|\hat{z}\|_{2}^{2}.$ (5)
@@
Consequently, injecting Gaussian noise in the Fourier domain preserves the $\ell_{2}$-sensitivity required for $(\varepsilon,\delta)$-DP, since the transform is unitary and does not amplify vector norms.
@@
Spectral filtering. A diagonal mask $\Phi=\mathrm{diag}(\varphi_{0},\dots,\varphi_{d-1})$ defines a linear filter $\mathcal{G}_{\Phi}(z)=\mathcal{F}^{-1}\left(\Phi\,\hat{z}\right)=\frac{1}{d}\sum_{k=0}^{d-1}\varphi_{k}\hat{z}_{k}\,e^{2\pi ikn/d}$ where $\hat{z}=\mathcal{F}(z)$. Equivalently, in matrix and convolution form,
$\mathcal{G}_{\Phi}$ $=\mathcal{F}^{-1}\Phi\mathcal{F},$ (6)
$(\mathcal{G}_{\Phi}z)_{n}$ $=\sum_{m=0}^{d-1}h_{(n-m)\bmod d}\,z_{m},\quad h=\mathcal{F}^{-1}\varphi.$ (7)
By Parseval’s identity,
$\|\mathcal{G}_{\Phi}z\|_{2}^{2}=\tfrac{1}{d}\sum_{k=0}^{d-1}|\varphi_{k}|^{2}\,|\hat{z}_{k}|^{2}\leq(\max_{k}|\varphi_{k}|^{2})\,\|z\|_{2}^{2}.$ (8)
@@
$\varphi_{k}=\begin{cases}1,&k<k_{0},\\[3.0pt] 1-\rho,&k\geq k_{0}.\end{cases}$ (9)
@@
Let $\Phi_{\rho}=\mathrm{diag}(\varphi_{0},\dots,\varphi_{d-1})$, then the filtered version of a privatized gradient $g=\nabla f+w$ is $\hat{g}=\mathcal{G}_{\Phi_{\rho}}(g)=\mathcal{F}^{-1}(\Phi_{\rho}\mathcal{F}(g)).$ When $w\sim\mathcal{N}(0,\sigma^{2}I)$, the transformed noise $\hat{w}:=\mathcal{F}^{-1}(\Phi_{\rho}\mathcal{F}(w))$ is still zero-mean but now has reduced energy in the low-frequency components:
$\mathbb{E}[\|\hat{w}_{<k_{0}}\|_{2}^{2}]\ll\mathbb{E}[\|\hat{w}\|_{2}^{2}],$ (10)
@@ 3.2 Gradient Dynamics with High-Frequency Differential Privacy
$\nabla F(x_{t})$ $=\nabla F(x_{t-1}+d_{t-1})$ (11)
$=\nabla F(x_{t-1})+\nabla^{2}F(x_{t-1})\,d_{t-1}$ (12)
$\quad+\tfrac{1}{2}\int_{0}^{1}\nabla^{3}F\bigl{(}(1-z)x_{t-1}+zx_{t}\bigr{)}[d_{t-1}]^{\otimes 2}\,dz,$ (13)
where $\boldsymbol{H}_{t}:=\nabla^{2}F(x_{t-1})\in\mathbb{R}^{d\times d}$ is approximated using privatized finite differences, and $d_{t-1}=x_{t}-x_{t-1}$. The observed gradient $g_{t}$ is a noisy, privatized estimate of the true gradient:
$g_{t}$ $=\frac{1}{B}\sum_{\xi\in\mathcal{B}_{t}}\text{clip}(\nabla f(x_{t},\xi),C)+w_{t}$ (14)
$=C_{t}\nabla F(x_{t})+w^{\prime}_{t},$ (15)
where $w^{\prime}_{t}$ contains both DP noise and subsampling noise, and $C_{t}$ is the effective observation operator with $\|C_{t}\|_{2}\leq 1$. Combining the update and observation equations:
@@
To enforce differential privacy while retaining useful structure, we first apply isotropic Gaussian noise $w_{t}\sim\mathcal{N}(0,\sigma_{w}^{2}I_{d})$ to the clipped gradient, followed by a deterministic frequency-domain transformation to shape the noise:
$g_{t}$ $=C_{t}\nabla F(x_{t})+w_{t}^{\prime},$ (16)
$w_{t}^{\prime}$ $=\mathcal{F}^{-1}\bigl{(}\Phi_{\rho}\odot\mathcal{F}(w_{t})\bigr{)},$ (17)
where $\Phi_{\rho}\in\mathbb{R}^{d}$ satisfies
$(\Phi_{\rho})_{k}$ $=\begin{cases}1,&0\leq k<k_{0},\\[3.0pt] 1-\rho e^{-\alpha(k-k_{0})},&k_{0}\leq k<d,\end{cases}$ (18)
with $k_{0}=\lfloor\lambda d\rfloor$, $\rho\in(0,1)$, and $\alpha>0$. This ensures that the privacy-preserving noise $w_{t}^{\prime}$ is spectrally shaped to occupy primarily high-frequency components, which contribute less to gradient descent, while preserving the $(\varepsilon,\delta)$-DP guarantee through post-processing. This approach facilitates improved recoverability of the informative low-frequency gradient content.